Which of the following statements are true? Please justify the answers.
1. a≥b⇔-a≤-b is an absolute inequality.
2. If A=φ,B={1,2},C={-1,-2}, then A×B×C has 4 elements.
3. The argument of 1+√(3)i is π/3.
4. A linear equation over R can have at most one root in C\R.
5. |x₁-x₂|=|x₁|-|x₂|∀x₁,x₂∈R
1
Expert's answer
2019-07-10T09:44:43-0400
  1. True
  2. False
  3. True
  4. False
  5. False

Justifying.

1)

"a \geq b, | \cdot(-1);"

"-a \leq -b"

2)

"A×B×C = \emptyset×B×C = \emptyset"

3)

"z=1+\sqrt{3}i" is a complex number;

"z=x+yi" is a complex number in the general form;

if "x > 0 \implies Arg(z) = \arctan(\frac y x) = \arctan(\frac {\sqrt{3}} {1}) = \arctan( \sqrt{3}) = \frac \pi 3"

4)

C\R set does not consist of elements of R set(real numbers), so a linear equation over R cannot have root over C\R.

5)

if x1 = -2 and x2 = 1, then

|x1 - x2| = | -2 - 1 | = | -3| = 3

|x1| - |x2| = |-2| - |1| = 2 - 1 =1,

therefore

|x1 - x2| "\neq" |x1| - |x2| "\forall" x1, x2 "\in R"